# Hybrid-pair superfluidity in a strongly driven Fermi gas

### v1

@@ B.2 Rabi-coupled three-component case
(1)   $-\epsilon_{-}\leq\bar{\xi}_{\bf k}$ (2)    $-\epsilon_{+}<\bar{\xi}_{\bf k}<-\epsilon_{-}$ (3)   $\bar{\xi}_{\bf k}<-\epsilon_{+}$
$E_{{\bf k}\uparrow}|_{\Delta_{1,2}=0}$ $\xi_{{\bf k}3}$ $-(\xi_{{\bf k}}+\epsilon_{-})$ $-(\xi_{{\bf k}}+\epsilon_{-})$
$E_{{\bf k}+}|_{\Delta_{1,2}=0}$ $\xi_{{\bf k}}+\epsilon_{+}$ $\xi_{{\bf k}}+\epsilon_{+}$ $-\xi_{{\bf k}3}$
$E_{{\bf k}-}|_{\Delta_{1,2}=0}$ $\xi_{{\bf k}}+\epsilon_{-}$ $-\xi_{{\bf k}3}$ $\xi_{{\bf k}}+\epsilon_{+}$
$\left(\frac{\partial^{2}E_{{\bf k}\uparrow}}{\partial\Delta_{1}^{2}}+R\frac{\partial^{2}E_{{\bf k}\uparrow}}{\partial\Delta_{1}\partial\Delta_{2}}\right)_{\Delta_{1,2}=0}$ $2\frac{c^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}+2\frac{csR+s^{2}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $-2\frac{c^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}$ $-2\frac{c^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}$
$\left(\frac{\partial^{2}E_{{\bf k}+}}{\partial\Delta_{1}^{2}}+R\frac{\partial^{2}E_{{\bf k}+}}{\partial\Delta_{1}\partial\Delta_{2}}\right)_{\Delta_{1,2}=0}$ $2\frac{s^{2}+csR}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $2\frac{csR+s^{2}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $-2\frac{c^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}-2\frac{s^{2}+csR}{\bar{\xi}_{\bf k}+\epsilon_{+}}$
$\left(\frac{\partial^{2}E_{{\bf k}-}}{\partial\Delta_{1}^{2}}+R\frac{\partial^{2}E_{{\bf k}-}}{\partial\Delta_{1}\partial\Delta_{2}}\right)_{\Delta_{1,2}=0}$ $2\frac{c^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}$ $-2\frac{c^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}-2\frac{s^{2}+csR}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $2\frac{s^{2}+csR}{\bar{\xi}_{\bf k}+\epsilon_{+}}$
$\left(\frac{\partial^{2}E_{{\bf k}\uparrow}}{\partial\Delta_{2}^{2}}+R^{-1}\frac{\partial^{2}E_{{\bf k}\uparrow}}{\partial\Delta_{1}\partial\Delta_{2}}\right)_{\Delta_{1,2}=0}$ $2\frac{s^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}+2\frac{c^{2}+R^{-1}cs}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $-2\frac{s^{2}-csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{-}}$ $-2\frac{s^{2}-csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{-}}$
$\left(\frac{\partial^{2}E_{{\bf k}+}}{\partial\Delta_{2}^{2}}+R^{-1}\frac{\partial^{2}E_{{\bf k}+}}{\partial\Delta_{1}\partial\Delta_{2}}\right)_{\Delta_{1,2}=0}$ $2\frac{c^{2}+csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $2\frac{c^{2}+csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $-2\frac{s^{2}-csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{-}}-2\frac{c^{2}+csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$
$\left(\frac{\partial^{2}E_{{\bf k}-}}{\partial\Delta_{2}^{2}}+R^{-1}\frac{\partial^{2}E_{{\bf k}-}}{\partial\Delta_{1}\partial\Delta_{2}}\right)_{\Delta_{1,2}=0}$ $2\frac{s^{2}-csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{-}}$ $-2\frac{s^{2}-csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{-}}-2\frac{c^{2}+csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $2\frac{c^{2}+csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$
@@
Due to the presence of the Fermi distributions in $A$ and $B$, we need to evaluate their integrands in the three regions separately. The results are summarized in Table 1. We see that, similarly to the quasiparticle energies, the partial derivatives can also be expressed in term of the non-interacting energies and coefficients. Using the property of the Fermi-Dirac distribution $f(x)=1-f(-x)$, we find
@@ Appendix C Analytical calculation of the order parameter in the BCS limit
[…]
@@
[…]
@@
[…]

### v2

@@ B.2 Rabi-coupled three-component case
(1) $\epsilon_{-}\leq\bar{\xi}_{\bf k}$ (2) $\epsilon_{+}<\bar{\xi}_{\bf k}<-\epsilon_{-}$ (3) $\bar{\xi}_{\bf k}<-\epsilon_{+}$
$E_{{\bf k}\uparrow}|_{\Delta_{1,2}=0}$ $\xi_{{\bf k}3}$ $(\xi_{{\bf k}}+\epsilon_{-})$ $(\xi_{{\bf k}}+\epsilon_{-})$
$E_{{\bf k}+}|_{\Delta_{1,2}=0}$ $\xi_{{\bf k}}+\epsilon_{+}$ $\xi_{{\bf k}}+\epsilon_{+}$ $\xi_{{\bf k}3}$
$E_{{\bf k}-}|_{\Delta_{1,2}=0}$ $\xi_{{\bf k}}+\epsilon_{-}$ $\xi_{{\bf k}3}$ $\xi_{{\bf k}}+\epsilon_{+}$
$\left(\frac{\partial^{2}E_{{\bf k}\uparrow}}{\partial\Delta_{1}^{2}}+R\frac{\partial^{2}E_{{\bf k}\uparrow}}{\partial\Delta_{1}\partial\Delta_{2}}\right)_{\Delta_{1,2}=0}$ $2\frac{c^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}+2\frac{csR+s^{2}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $2\frac{c^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}$ $2\frac{c^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}$
$\left(\frac{\partial^{2}E_{{\bf k}+}}{\partial\Delta_{1}^{2}}+R\frac{\partial^{2}E_{{\bf k}+}}{\partial\Delta_{1}\partial\Delta_{2}}\right)_{\Delta_{1,2}=0}$ $2\frac{s^{2}+csR}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $2\frac{csR+s^{2}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $2\frac{c^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}-2\frac{s^{2}+csR}{\bar{\xi}_{\bf k}+\epsilon_{+}}$
$\left(\frac{\partial^{2}E_{{\bf k}-}}{\partial\Delta_{1}^{2}}+R\frac{\partial^{2}E_{{\bf k}-}}{\partial\Delta_{1}\partial\Delta_{2}}\right)_{\Delta_{1,2}=0}$ $2\frac{c^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}$ $2\frac{c^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}-2\frac{s^{2}+csR}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $2\frac{s^{2}+csR}{\bar{\xi}_{\bf k}+\epsilon_{+}}$
$\left(\frac{\partial^{2}E_{{\bf k}\uparrow}}{\partial\Delta_{2}^{2}}+R^{-1}\frac{\partial^{2}E_{{\bf k}\uparrow}}{\partial\Delta_{1}\partial\Delta_{2}}\right)_{\Delta_{1,2}=0}$ $2\frac{s^{2}-csR}{\bar{\xi}_{\bf k}+\epsilon_{-}}+2\frac{c^{2}+R^{-1}cs}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $2\frac{s^{2}-csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{-}}$ $2\frac{s^{2}-csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{-}}$
$\left(\frac{\partial^{2}E_{{\bf k}+}}{\partial\Delta_{2}^{2}}+R^{-1}\frac{\partial^{2}E_{{\bf k}+}}{\partial\Delta_{1}\partial\Delta_{2}}\right)_{\Delta_{1,2}=0}$ $2\frac{c^{2}+csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $2\frac{c^{2}+csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $2\frac{s^{2}-csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{-}}-2\frac{c^{2}+csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$
$\left(\frac{\partial^{2}E_{{\bf k}-}}{\partial\Delta_{2}^{2}}+R^{-1}\frac{\partial^{2}E_{{\bf k}-}}{\partial\Delta_{1}\partial\Delta_{2}}\right)_{\Delta_{1,2}=0}$ $2\frac{s^{2}-csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{-}}$ $2\frac{s^{2}-csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{-}}-2\frac{c^{2}+csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$ $2\frac{c^{2}+csR^{-1}}{\bar{\xi}_{\bf k}+\epsilon_{+}}$
@@
Due to the presence of the Fermi distributions in $A$ and $B$, we need to evaluate their integrands in the three regions separately. The results are summarized in Table 1. We see that, similarly to the quasiparticle energies, the partial derivatives can also be expressed in term of the non-interacting energies and coefficients. Using the property of the Fermi-Dirac distribution f(x)=1f(x)f(x)=1-f(-x), we find
@@ Appendix C Analytical calculation of the order parameter in the BCS limit
$\frac{\pi}{k_{F}a_{13}}=$ $\underbrace{\int_{0}^{1}dx\sqrt{x}\left\{-\frac{c^{2}}{\sqrt{\zeta_{\bf x}^{2}/4+|\Delta/E_{F}|^{2}c^{2}}}+\frac{1}{x}\right\}}_{I_{1}}$
$+\underbrace{\int_{1}^{\infty}dx\sqrt{x}\left\{-\frac{c_{\bf x}^{2}}{\sqrt{\zeta_{\bf x}^{2}/4+|\Delta/E_{F}|^{2}c_{\bf x}^{2}}}+\frac{1}{x}\right\}}_{I_{2}},$ (91)
@@
$I_{2}$ $\simeq\underbrace{\int_{1}^{\infty}dx\left(\sqrt{x}-1\right)\left\{-\frac{2c_{\bf x}^{2}}{\zeta_{\bf x}}+\frac{1}{x}\right\}}_{I_{2}^{a}}$
$+\underbrace{\int_{1}^{\infty}dx\left\{-\frac{c_{\bf x}^{2}}{\sqrt{\zeta_{\bf x}^{2}/4+|\Delta/E_{F}|^{2}c_{\bf x}^{2}}}+\frac{1}{x}\right\}}_{I_{2}^{b}},$ (93)
@@
$I_{2}^{a}=\int_{1}^{\infty}dx\left(\sqrt{x}-1\right)\left\{-\frac{2c^{2}}{\bar{\xi}_{\bf x}+\epsilon_{-}}-\frac{2s^{2}}{\bar{\xi}_{\bf x}+\epsilon_{+}}+\frac{1}{x}\right\},$ (94)